OpenStudy (anonymous):

Solve the system of equations using the substitution method. 1. y=x^2+2x-1 y=-2x-4 2. y=2x^2-5 y=11-4x

OpenStudy (solomonzelman):

for the first problem. you are given (in the second equation) that \(\large\color{black}{ \displaystyle y }\) is same thing as \(\large\color{black}{ \displaystyle -2x-4 }\) So, all you need to do is to substitute this (the -2x-4 ) instead of y into the first equation.

OpenStudy (anonymous):

oh ok thank you that makes more sense

OpenStudy (solomonzelman):

the substitution will occur the following way: YOUR FIRST EQUATION \(\large\color{black}{ \displaystyle \color{red}{y}=x^2+2x-1 }\) SUBSTITUTION \(\large\color{black}{ \displaystyle \color{red}{-2x-4}=x^2+2x-1 }\)

OpenStudy (solomonzelman):

(I am lagging, excuse me)

OpenStudy (solomonzelman):

simplify and factor.